HelpMePlz:

Use a calculator to find the Standard Deviation and the mean for the following set of data: 32, 61, 14, 23, 5, 25, 38

Champion:

Yes, mean is \(approximately\) equal to 28.3. To be exact, it is 28+2/7=28 2/7 If the numbers form a population, the mean is denoted \(\mu\). If the numbers form a sample from a population, the mean is denoted xbar. There are two formulas for standard deviation. One is for a population, \(\sigma\) The other is for a sample taken from a population, and so it is an approximation, s. \(\sigma=\sqrt{\sum(x-\mu)^2/N}\) where N is the size of the population \(s=\sqrt{\sum(x-xbar)^2/(n-1)}\) where n is the size of the sample. If you still have doubts, read the following: http://davidmlane.com/hyperstat/A16252.html